OpenStudy (anonymous):

i just want someone to double check me so that i know i am correct i believe the answer is equation 1 has less solutions than equation2 Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 -5|3x - 2| = -15 |2x - 9| = 11

OpenStudy (anonymous):

\[-5\left| 3x-2 \right|=-15\] \[\left| 3x-2 \right|=3\] \[3x-2=3\] \[3x=5\] \[x=\frac{ 5 }{ 3 }\] or \[3x-2=-3\] \[3x=-1\] \[x=-\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

\[\left| 2x-9 \right|=11\] \[2x-9=11\] \[2x=20\] \[x=10\] or \[2x-9=-11\] \[2x=-2\] \[x=-1\]

OpenStudy (anonymous):

both equations have two solutions

OpenStudy (anonymous):

ooh so i was wrong , i knew something was up with my answer thank yo @le0n

OpenStudy (anonymous):

you're welcome